OpenStudy (anonymous):

find how much time passes before the Blood Alcohol level reaches .04, we made C(t)=.04, and then solved the original equation (not the derivative) for t. This is the work that I've done for the problem, but I know it's not right, the right answer is t=5.66 .04=(.12t)(e^(-t/2) ln(.04)=ln((.12t)(e^(-t/2)) ln(.04)=(.12t)(-t/2) 2ln(.04)=(.12t)(-t) 2ln(.04)=-.12t^2 (2ln(.04))/-.12=t^2 Sqrt(2ln(.04))/-.12) = t

OpenStudy (anonymous):

What is the original equation? C(t) = C(final)[1 - exp(-kt)] ? Or ?

OpenStudy (anonymous):

C (t) = 0.12te^(-t/2) and c'(t)=.06(t-2)e^(-t/2)

OpenStudy (anonymous):

0.04 = 0.12 t exp(-t/2) ln(0.04) = ln(0.12) + ln(t) - t/2 solve for t. t = 2[ln(3t)] Plot y = t and y = 2[ln(3t)] vs. t and see value of t where they intersect.

OpenStudy (anonymous):

How did you simplify to t=2[ln(3t)] for the problem that this is a part of, I need to show all of my steps for solving of t

OpenStudy (anonymous):

ln(0.12) - ln(0.04) = ln(0.12/0.04) = ln(3) ln(3) + ln(t) = ln(3t)

OpenStudy (anonymous):

Thank you so so so much!